OpenStudy (anonymous):

the perimeter of a rectangle is twice the sum of its length and its width. the perimeter is 22 meters and its length is 2 meters more than twice its width. what is its length? HELP

OpenStudy (mathmath333):

\(2(l+w)=22\\l=2w+2\\\)

OpenStudy (anonymous):

P=2(l+w) Solve for l 22=2((2+2w)+w 4+4w+w=22 4+5w=22 22-4=18 18/5=3.6 l=3.6

OpenStudy (anonymous):

And, to check that: Solve for w 22=2(3.6+w) 22=7.2+2w 22-7.2=14.8 14.8/2=7.4 w=7.4 Check 22=2(3.6+7.4) 22+2(11) 22=22